Hi there, I really need some help with solving word problems! Here is the question:
The sum of Louise's age and Todd's age is 34 years. Five years ago, the sum of twice Louise's age and three times Todd's age was 61 years. How old are they now?

Answer by arunpaul(104)

Let Louise's age =x
and Todd's age =y
as per question x+y = 34 -----eq(1)
five years ago Louise's age was = x-5
and Todd's age was =y-5
and 2(x-5)+3(y-5)=61
or 2x-10+3y-15=61
or 2x+3y-25=61
or 2x+3y=61+25=86-------eq(2)
solving eq(1) and eq(2)we have
i.e eq(1)-eq(2)
2x+2y=68
-2x-3y=-86
-----------
-y =-18
or y =18
so x = 34-18=16
so Todd's age =18 years
and Louise's age =16 years